9. In Linumusitatissimum (flax, used to make linen textile), the blue color of the flowers (b) and the long fiber length (l) are recessive traits. A farmer mate true breeding plants with short fibers and white flowers with plants with blue flowers and long fibers.
a) What are the genotypes of these plants?
b) What is the geno type of the F1generation?
c) If plants from F1 are auto-fertilized and 800 F2 plants are obtained,
1. calculate the expected number of plants with white flowers and long fibers
2. determine the percentage of the different genotypes among these (white flowers and long fibers).

Could you please give a detailed explannation on this exercise? Thank you!
Answer #1

Blue coloured flowers = bb

White coloured flowers = BB, Bb

Long fibers = ll

Short fibers = LL, Ll

A.

True breeding plant with short fibers and white flowers = LLBB

True breeding plant with long fibers and blue flowers = llbb

B.

LLBB × llbb = F1 generation = LlBb (short fibers, white flowers)

C.

LlBb × LlBb = F2 generation

Gametes LB Lb lB lb
LB LLBB (short, white) LLBb (short, white) LlBB (short, white) LlBb (short, white)
Lb LLBb (short, white) LLbb (short, blue) LlBb (short, white) Llbb (short, blue)
lB LlBB (short, white) LlBb (short, white) llBB (long, white) llBb (long, white)
lb LlBb (short, white) Llbb (short, blue) llBb (long, white) llbb (long, blue)

Expected Phenotypic ratio = 9:3:3:1

9/16 = L_B_ (short, white)

3/16 = llB_ (long, white)

3/16 = L_bb (short, blue)

1/16 = llbb (long, blue)

Expected Genotypic ratio = 1:1:2:2:4:2:2:1:1

1/16 = LLBB

1/16 = LLbb

2/16 = LLBb

2/16 = LlBB

4/16 = LlBb

2/16 = Llbb

2/16 = llBb

1/16 = llBB

1/16 = llbb

1. Expected number of plants with white flowers and long fibers = (3/16) × 800 = 150

2. Proportion of genotype of white flowers and long fibers (llB_)

llBb = 2/16

Number of plants with llBb = (2/16) × 800 = 100

llBB = 1/16

Number of plants with llBB = (1/16) × 800 = 50

Please rate.